OpenStudy (anonymous):

what is the simplified form of 15x^5/24y^8 divided by 4x^2/8y^4?

OpenStudy (anonymous):

Use mathway to answer your question

OpenStudy (anonymous):

@ysiad i tried but i cant type it out right

OpenStudy (solomonzelman):

\[\huge\color{blue}{ \frac{15x^5}{24y^8} \div \frac{4x^2}{8y^4}~~~->~~~~ } \frac{15x^5}{24y^8} \times \frac{8y^4}{4x^2} \] can you do the problem in black?

OpenStudy (anonymous):

i simplified my answer and got 5x^3/4y^4

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

thank you

OpenStudy (solomonzelman):

Correct!

OpenStudy (solomonzelman):

I though you were than thought you weren't, and then re-looked at it, and yes you are correct!

OpenStudy (anonymous):

lol ok thanks

OpenStudy (solomonzelman):

Yw!